find the equation of the line containing (-3,1)and perpendicular to the line 2x+3y=5

Answer by tutorcecilia(2152)

The slopes of perpendicular lines are are negative reciprocals of each other.
.
Line #1:
2x+3y=5 [Re-write in the slope-intercept form by solving for "y"]
y=-2/3x+5/3 [The slope (m) = -2/3]
.
Line 2 has points (-3,1). The slope is the negative reciprocal of (-2/3) = (-(-3/2) = 3/2
y=mx+b [Plug in the pionts and slope for Line 2]
1=(3/2)(-3)+b [Solve for b]
1=-9/2+b
1+9/2=-9/2 - 9/2+b
18/2=b
9=b
y=3/2x+9 [The equation of Line #2 that is perpendicular to Line #1]